# Tropical Gradient Descent

### v1

@@ 3.2 Convergence Guarantees
In this section we prove the theoretical guarantees of tropical descent for max-tropical location problems. We show that tropical descent must converge to the tropical convex hull of the data, and hence will necessarily find a minimum of a $\triangle_{\max}$-star-convex problem. Throughout this section, we assume the step sizes $a_{n}$ satisfy $a_{n}\rightarrow 0,\sum a_{n}=\infty$.
We first highlight the key difference in stability between classical descent and tropical descent; if some component of the gradient is locally zero, tropical descent cannot converge in that neighbourhood.
@@
Assume that for all $n$, $\nabla f(\mathbf{t}_{n})\neq\mathbf{0}$. Suppose in some bounded open $U$, $\nabla f_{i}$ is identically 0. Then no point in $\operatorname{int}U$ is a convergence point for $\mathbf{t}_{n}$; in particular, if $\mathbf{t}_{m}\in U$ we will eventually leave $U$.
Proof of Proposition 24.
Assume otherwise. Then for all $n\geq m$, we have $\nabla f(\mathbf{t}_{n})_{i}=0$, so $t_{ni}$ is constant. At each step, we decrease at least one coordinate by $a_{n}$, so as $n\rightarrow\infty$:
$\sum_{j}t_{nj}\leq\sum_{j}t_{mj}-\sum_{n\geq m}a_{n}\rightarrow-\infty.$
Therefore $\|\mathbf{t}_{n}\|_{{\rm tr}}$ will become arbitrarily large, as $\min_{j}t_{nj}\rightarrow-\infty$ while $\max_{j}t_{nj}\geq t_{ni}$. Then we must have left $U$; contradiction. ∎
As a result, local minima such as the valleys in figures 4 and 3 can be stable with respect to classical descent but unstable with respect to tropical descent. We next prove an opposing result; tropical gradient descent must converge to the tropical convex hull of the data points. This result stems from the following observation on the gradient of $\triangle_{\max}$-star-quasi-convex functions.
Remark 25.
We note that for a $\triangle_{\max}$-star-quasi-convex function with kernel $\mathbf{v}$, we have that
$\nabla f(\mathbf{x})_{j}>0\Rightarrow j\in\operatorname{argmax}_{i}(x_{i}-v_{i}).$
@@
Suppose $f$ is a max-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Let $V=\operatorname{tconv}_{\min}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta_{i}(\mathbf{t})=\sum_{k\in[K]}\sum_{j\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
For any $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}\leq\epsilon/N$. Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$. Then almost surely, for all $m\geq M_{2}$ we have:
@@
Proof.
For each $i\leq N$, we define the functions
$\delta_{i}(\mathbf{t})=\min_{k\in[K]}[\sum_{j}[t_{j}-v_{kj}-t_{i}+v_{ki}]^{+}].$
@@
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\max_{j}\delta_{j}(\mathbf{t}_{m})$.
Proof.
As discussed in Maclagan & Sturmfels, (2015), we have the projection map onto $V$ given by:
$\pi_{V}(\mathbf{t})=\lambda_{1}\odot\mathbf{x}_{1}\oplus\dots\oplus\lambda_{K}\odot\mathbf{x}_{K},\qquad\text{ where }\lambda_{k}=\max_{i}(t_{i}-x_{ki})$
So the distance to $V$ is given by $d(\mathbf{t},\pi_{V}(\mathbf{t}))$, and we have:
@@
$=\|(\min_{k}(\max_{i}(t_{i}-x_{ki})+x_{kj}-t_{j}))_{j}\|$
$\leq\max_{j}(\min_{k}(\max_{i}(t_{i}-x_{ki})+x_{kj}-t_{j}))$
$\leq\max_{j}\delta_{j}(\mathbf{t})$
@@
Proof.
For all $k\in[K],j\in[N]$, we have that $t_{j}-v_{kj}-t_{i}+v_{ki}$ is $1$-Lipschitz in $\mathbf{t}$. Therefore $\delta_{i}$ is $N$- Lipschitz in $\mathbf{t}$, and if $\delta_{i}(\mathbf{t}_{m})=0,\delta_{i}(\mathbf{t}_{m+1})\leq Na_{m}<\epsilon$. ∎
@@
If $\delta_{i}(\mathbf{t}_{m})>0$ then $\delta_{i}(\mathbf{t}_{m+1})\leq\delta_{i}(\mathbf{t}_{m})$.
Proof.
Suppose $\delta_{i}(\mathbf{t}_{m})>0$. Then $\forall k$, $\max_{j}(t_{j}-v_{kj})>t_{i}-v_{ki}$. By remark 25, $\nabla f_{i}\leq 0$ and $d_{mi}=0$. Therefore $\delta_{i}(\mathbf{t}_{m+1})\leq\delta_{i}(\mathbf{t}_{m})$. ∎
@@
Let $M$ be such that $s_{M}\geq\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$. Then for some $M_{1}\leq m\leq M$, $\delta_{i}(\mathbf{t}_{m})<\epsilon$.
Proof.
We note that $K\delta_{i}(\mathbf{t})\leq\Delta_{i}(\mathbf{t})$; hence if $\Delta_{i}(\mathbf{t}_{M_{1}})<K\epsilon$ then we are done.
We now assume $\Delta_{i}(\mathbf{t}_{M_{1}})\geq K\epsilon$ and prove the result by contradiction; suppose $\delta_{i}(\mathbf{t}_{m})>\epsilon$ for all $M_{1}\leq m\leq M$. Therefore at each step we have $\nabla f_{i}\leq 0$, $d_{mi}=0$ and $t_{mi}$ is fixed for $n\leq M$.
As $\nabla f\neq\mathbf{0}$, there is some $j$ such that $\nabla f_{j}>0$, and let $k$ be such that $j=\operatorname{argmax}_{l}t_{ml}-v_{kl}$. As $\delta_{i}(\mathbf{t}_{m})>\epsilon$:
$t_{j}-v_{kj}-t_{i}+v_{ki}>\epsilon/N\geq a_{m}.$
Hence this term in the summation of $\Delta_{i}$ decreases by $a_{m}$, while other terms are non-increasing. We conclude that $\Delta_{i}(\mathbf{t}_{M})\leq\Delta_{i}(\mathbf{t}_{M_{1}})-s_{M}+s_{M_{1}}\leq K\epsilon$. Contradiction. ∎
Finally, from claims 28 to 30, we conclude that $\forall m\geq M_{2},i\leq N$, we have $\delta_{i}(\mathbf{t}_{m})<\epsilon$, and by claim 27, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$. ∎
As proven in Comăneci, (2023), tropical location problems have minima in the tropical convex hull of the data. This result says that according to tropical gradient descent, only such minima can be stable. As a corollary, we can achieve the same result for min-tropical location problems by using negative tropical descent directions.
@@
Suppose $f$ is a min-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined via negative tropical descent directions, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Let $V=\operatorname{tconv}_{\max}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta^{\prime}_{j}(\mathbf{t})=\sum_{k\in[K]}\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
For any initial $\mathbf{t}_{0}$ and $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}\leq\epsilon/N$. Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta^{\prime}_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$. Then almost surely, for all $m\geq M_{2}$ we have:
@@
Despite following from the previous results, the following corollary may be considered the most powerful result of this paper; tropical descent will necessarily minimise any max-tropically quasi-convex function.
Corollary 32.
Let $f(\mathbf{t})$ be a max-tropically quasiconvex function, and let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\leq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$.
We conclude by establishing the convergence rates of tropical descent in practice; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume gradients have bounded norm as in the the case of piecewise linear objective functions. In this case, we prove convergence at a rate of $O(\sqrt{m})$.
@@
Suppose $f$ is a max-tropical location problem as in Theorem 26. Assume that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall k\leq K,n\in\mathbb{N}:\,d_{{\rm tr}}(\mathbf{t}_{n},\mathbf{x}_{k})\leq D$. We define $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$. Then for all $m$ satisfying
$\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}>\sqrt{2},$
@@
In particular, if $f$ is max-tropically quasiconvex, then there is some global minimum $\mathbf{t}^{*}$ such that
@@
Proof.
We fix $m$ and define:
$M_{1}=\left\lfloor\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}\right)^{2}\right\rfloor.$
@@
$\sqrt{M_{1}}$ $\leq\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1},$
$0\leq\sqrt{M_{1}}$ $\leq\frac{L}{2N}\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}+\sqrt{\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}\right)^{2}+\frac{4N^{2}(LKN-1)}{L^{2}}}\right).$
@@
$0\geq\frac{N}{L}M_{1}+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)\sqrt{M_{1}}+\frac{N-LKN^{2}}{L}$
$\frac{N\sqrt{M_{1}}\sqrt{M_{1}+1}}{L}\leq\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}\right)\sqrt{M_{1}}+KN^{2}$
$KDN(N-1)/2+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}$
@@
For all $n\geq M_{1}$, we have $a_{n}\leq\epsilon_{m}/N$.
Proof.
We have assumed that
$\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}>\sqrt{2}$
Then:
$M_{1}$ $\geq\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}\right)^{2}-1$
$\geq\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}-1\right)^{2}$
$\sqrt{M_{1}}$ $\geq\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}-1$
$\geq(\sqrt{m+1}-KDNL(N-1)/4\alpha)-1$
Hence:
$a_{n}$ $\leq\frac{2\alpha}{M_{1}^{1/2}}$
$\leq\frac{2\alpha}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}/N$
@@
For all $n\geq m$, we have $\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
Proof.
We note that
$\max_{i}\Delta_{i}(\mathbf{t})$ $\leq\sum_{i}\Delta_{i}(\mathbf{t})$
$=\sum_{k\leq K}\sum_{i\leq j\leq N}|t_{j}-x_{kj}-t_{i}+x_{ki}|$
$=\sum_{k\leq K}\sum_{i\leq j\leq N}d_{{\rm tr}}(\mathbf{t},\mathbf{x}_{k}),$
and hence
$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})$ $\leq\frac{KDN(N-1)}{2}.$
@@
$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq KDN(N-1)/2-2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
$\leq\frac{2\alpha KN}{\sqrt{M_{1}}}$
$\leq K\epsilon_{m}$
Hence by Theorem 26, we have almost surely that
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon_{m}$
The bound for the tropically quasiconvex case follows from Corollary 32.

### v2

@@ 3.2 Convergence Guarantees
In this section we prove the theoretical guarantees of tropical descent for min-tropical location problems. We show that tropical descent must converge to the tropical convex hull of the data, and hence will necessarily find a minimum of a $\triangle_{\min}$-star-convex problem. Throughout this section, we assume the step sizes $a_{n}$ satisfy $a_{n}\rightarrow 0,\sum a_{n}=\infty$.
We first highlight the key difference in stability between classical descent and tropical descent; if some component of the derivative is locally zero, tropical descent cannot converge in that neighbourhood.
@@
Let $\mathbf{t}_{n}$ be some sequence defined by tropical descent. Assume that for all $n$, $\nabla f(\mathbf{t}_{n})\neq\mathbf{0}$. Suppose in some bounded open $U$, $\nabla f_{i}$ is identically 0. Then no point in $\operatorname{int}U$ is a convergence point for $\mathbf{t}_{n}$; in particular, if $\mathbf{t}_{m}\in U$ there is some $n>m$ such that $\mathbf{t}_{n}\notin U$.
Proof Assume otherwise. Then for all $n\geq m$, we have $\nabla f(\mathbf{t}_{n})_{i}=0$, so $t_{ni}$ is constant. At each step, we increase at least one coordinate by $a_{n}$, so as $n\rightarrow\infty$:
$\sum_{j}t_{nj}\geq\sum_{j}t_{mj}+\sum_{n\geq m}a_{n}\rightarrow\infty.$
Therefore $\|\mathbf{t}_{n}\|_{{\rm tr}}$ will become arbitrarily large, as $\max_{j}t_{nj}\rightarrow\infty$ while $\min_{j}t_{nj}\leq t_{mi}$. Then we must have left $U$; contradiction. ∎
As a result of the proposition above, local minima such as the valleys in Figures 3 and 4 can be stable with respect to classical descent but unstable with respect to tropical descent.
We next prove the main convergence result of this work; tropical descent must converge to the tropical convex hull of the data points.
@@
Suppose $f$ is a min-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Let $V=\operatorname{tconv}_{\max}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta_{j}(\mathbf{t})=\sum_{k\in[K]}\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
For any $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}<\epsilon/N$. Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$, where $s_{m}$ is the partial sum of the step size sequence $a_{m}$. Then for all $m\geq M_{2}$ we have:
@@
Proof For each $i\leq N$, we define the functions
$\delta_{j}(\mathbf{t})=\min_{k\in[K]}\left[\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}\right].$
@@
For all $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, we have $d_{{\rm tr}}(\mathbf{t},V)\leq\max_{j}\delta_{j}(\mathbf{t})$.
Proof As discussed by Maclagan & Sturmfels, (2015), there is a well-defined metric projection map $\pi_{V}$ onto $V$ given by:
$\pi_{V}(\mathbf{t})=\lambda_{1}\odot\mathbf{x}_{1}\boxplus\dots\boxplus\lambda_{K}\odot\mathbf{x}_{K},$
[…]
The distance from $\mathbf{t}$ to $V$ is then given by $d(\mathbf{t},\pi_{V}(\mathbf{t}))$, and we have:
@@
$=\|(\max_{k}(\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj})_{j\leq N}\|.$
We note that for all $j,k$, the term $\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj}$ is negative. Hence:
$d(\mathbf{t},V)$ $\leq 0-\min_{j}(\max_{k}(\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj}))$
$\leq\max_{j}(\min_{k}(t_{j}-x_{kj}-\min_{i}(t_{i}-x_{ki})))$
$\leq\max_{j}\delta_{j}(\mathbf{t}).$
The claim is proved. ∎
We will now prove that for all $m\geq M_{2}$, we have that $\delta_{j}(\mathbf{t}_{m})\leq\epsilon$.
@@
Proof For all $k\in[K],j\in[N]$, we have that $t_{mj}-x_{kj}-t_{mi}+x_{ki}$ is $1$-Lipschitz in $\mathbf{t}$ (with respect to $d_{\mathrm{tr}}$). Therefore $\delta_{j}$ is $N$-Lipschitz in $\mathbf{t}$. By the tropical descent rule:
$\delta_{j}(t_{m+1})$ $\leq\delta_{j}(\mathbf{t}_{m})+Nd_{\mathrm{tr}}(\mathbf{t}_{m},\mathbf{t}_{m+1})$
$=0+Na_{m}$
$<\epsilon.$
The claim is proved. ∎
@@
If $\delta_{j}(\mathbf{t}_{m})>0$ then $\delta_{j}(\mathbf{t}_{m+1})\leq\delta_{j}(\mathbf{t}_{m})$.
Proof Suppose $\delta_{j}(\mathbf{t}_{m})>0$. Then $\forall k$, $t_{mj}-x_{kj}>\min_{i}(t_{mi}-x_{ki})$, so by Lemma 25 we have $\nabla f_{j}\geq 0$. The tropical steepest descent direction therefore has $d_{mj}=0$. We conclude that
$\forall\,k,i:\quad$ $t_{(m+1)j}-x_{kj}-t_{(m+1)i}+x_{ki}\leq t_{mj}-x_{kj}-t_{mi}+x_{ki},$
$\Rightarrow\quad\delta_{j}(\mathbf{t}_{m+1})\leq\delta_{j}(\mathbf{t}_{m}).$
The claim is proved. ∎
@@
For some $M_{1}\leq m\leq M_{2}$, $\delta_{j}(\mathbf{t}_{m})<\epsilon$.
Proof We note that $K\delta_{j}(\mathbf{t})\leq\Delta_{j}(\mathbf{t})$; hence if $\Delta_{j}(\mathbf{t}_{M_{1}})<K\epsilon$ then we are done.
We assume $\delta_{j}(\mathbf{t}_{m})>\epsilon$ for all $M_{1}\leq m\leq M_{2}$, and prove the result by contradiction. Then as in the proof of the previous claim, at each step we have $\nabla f_{j}\geq 0$, $d_{mj}=0$ and $t_{mj}$ is fixed for $M_{1}\leq m\leq M_{2}$.
We have assumed that for each $m$, $\nabla f\neq\mathbf{0}$ and so there is some $i$ such that $\nabla f_{i}<0$, and by Lemma 25, there is some $k$ such that $i\in\text{argmin}_{\ell}t_{m\ell}-v_{k\ell}$. Hence:
$\forall\,\ell:\quad t_{mj}-v_{kj}-t_{mi}+v_{ki}$ $\geq t_{mj}-v_{kj}-t_{m\ell}+v_{k\ell},$
$\Rightarrow\quad t_{mj}-v_{kj}-t_{mi}+v_{ki}$ $\geq\delta_{j}(\mathbf{t}_{m})/N,$
$>\epsilon/N,$
$\geq a_{m}.$
Hence this term in the summation of $\Delta_{j}(\mathbf{t}_{m})$ decreases by $a_{m}$, while other terms are non-increasing as $t_{mj}$ is fixed.
Noting that $K\delta_{j}(\mathbf{t})\leq\Delta_{j}(\mathbf{t})$, we conclude
$K\delta_{j}(\mathbf{t}_{M_{2}})$ $\leq\Delta_{j}(\mathbf{t}_{M_{2}}),$
$\leq\Delta_{j}(\mathbf{t}_{M_{1}})-s_{M_{2}}+s_{M_{1}},$
$\leq K\epsilon.$
Contradiction. So the claim is proved. ∎
Finally, from the claims above, we conclude that $\forall m\geq M_{2},i\in[N]$, we have $\delta_{i}(\mathbf{t}_{m})<\epsilon$, and so by our first claim, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$. ∎
As proven by Comăneci, (2024), tropical location problems have minima in the tropical convex hull of the data. The theorem above says that according to tropical descent, only such minima can be stable. We cannot do much better in terms of convergence results for general location problems; they can have disconnected sublevel sets so we would not expect gradient methods to find a global minimum. We would require some degree of convexity to guarantee global solvability, which is exactly what the following corollary gives us; tropical descent will necessarily converge to a minimum of any $\triangle_{\min}$-star-quasi-convex function.
@@
Let $f(\mathbf{t})$ be a $\triangle_{\min}$-star-quasi-convex function, and let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\leq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$ where $M_{2}$ is as in Theorem 29.
As a further corollary, we note that we can achieve the same result for max-tropical location problems by using max-tropical descent directions.
Corollary 35.
Suppose $f$ is a max-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined via max-tropical descent directions, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Let $V=\operatorname{tconv}_{\min}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta^{\prime}_{i}(\mathbf{t})=\sum_{k\in[K]}\sum_{j\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
For any initial $\mathbf{t}_{0}$ and $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}\leq\epsilon/N$. Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta^{\prime}_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$. Then for all $m\geq M_{2}$ we have:
@@
Remark 36.
The results above indicate convergence to $V$ or $\mathbf{t}^{*}$ in the tropical topology of $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, but $\mathbf{t}_{n}$ cannot not converge in the Euclidean topology of $\mathbb{R}^{N}$ due to the divergent partial sums of $\alpha_{n}$.
We conclude by establishing the convergence rate of tropical descent in practice; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume derivatives have bounded tropical norm as is the case for piecewise linear 1-Lipschitz objective functions $h_{k}$. In this case, we prove convergence at a rate of $O(\sqrt{m})$.
@@
Suppose $f$ is a min-tropical location problem as in Theorem 29. Assume that there is some $L\in\mathbb{N}$ such that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall k\leq K,n\in\mathbb{N}:\,d_{{\rm tr}}(\mathbf{t}_{n},\mathbf{x}_{k})\leq D$. We define $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$. Then for all $m$ satisfying
$1<\xi(m)\coloneqq\frac{\sqrt{m+1}}{2}-\frac{KDNL(N-1)}{8\alpha}+\sqrt{\left(\frac{\sqrt{m+1}}{2}-\frac{KDNL(N-1)}{8\alpha}\right)^{2}+LKN-1},$
@@
In particular, if $f$ is a $\triangle_{\min}$-star-quasi-convex function, then there is some global minimum $\mathbf{t}^{*}$ such that
@@
Proof We fix $m$ and define:
$M_{1}=\left\lfloor\xi(m)^{2}\right\rfloor.$
@@
$0\leq\sqrt{M_{1}}\leq\xi(m)$
@@
$\frac{N}{L}(M_{1}+1)+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)\sqrt{M_{1}}-KN^{2}\leq 0,$
$\frac{N\sqrt{M_{1}}\sqrt{M_{1}+1}}{L}+\frac{KDN^{2}(N-1)\sqrt{M_{1}}}{4\alpha}-\frac{N\sqrt{M_{1}}\sqrt{m+1}}{L}\leq KN^{2},$
$\frac{KDN(N-1)}{2}+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}.$
@@
For all $n\geq M_{1}$ and $m$ such that $1<\xi(m)$, we have $a_{n}\leq\epsilon_{m}/N$.
Proof We have assumed that $1<\xi(m)$, so:
$M_{1}$ $\geq\xi(m)^{2}-1,$
$\geq\left(\xi(m)-1\right)^{2},$
$\sqrt{M_{1}}$ $\geq\xi(m)-1,\geq(\sqrt{m+1}-KDNL(N-1)/4\alpha)-1.$
Hence, by the step size definition:
$a_{n}$ $\leq\frac{2\alpha}{\sqrt{M_{1}}},$
$\leq\frac{2\alpha}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}/N.$
The claim is proved. ∎
@@
For all $n\geq m$, we have $\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
Proof We note that
$\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})$ $\leq\sum_{j}\Delta_{j}(\mathbf{t}_{M_{1}}),$
$=\sum_{k\leq K}\sum_{i\leq j\leq N}|t_{M_{1}j}-x_{kj}-t_{M_{1}i}+x_{ki}|,$
$\leq\sum_{k\leq K}\sum_{i\leq j\leq N}d_{{\rm tr}}(\mathbf{t}_{M_{1}},\mathbf{x}_{k}),$
$\leq\frac{KDN(N-1)}{2}.$
@@
$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq KDN(N-1)/2+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1}),$
$\leq\frac{2\alpha KN}{\sqrt{M_{1}}},$
$\leq K\epsilon_{m}.$
Hence by Theorem 29, we have that
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon_{m}.$
The bound for the tropically quasi-convex case follows from Corollary 34. ∎